OpenStudy (anonymous):

y = x^2 - 4x + 1 into vertex form

OpenStudy (anonymous):

\[ y = x^2 -4 x + 4 -4 +1 \] Can you finish it?

OpenStudy (anonymous):

\[ y = x^2 -4 x + 4 -4 +1\\ y=(x-2)^2 -3 \]

OpenStudy (anonymous):

would it give me coordinates of, (2,3) ?